# Iterative Partition Search Variational Quantum Algorithm for Solving Shortest Vector Problem

### v1

@@ IV.3  Comparison with Existing Algorithms
We now compare the effectiveness of IPSA against existing algorithms, namely the various $k$-PSA configurations and IQOAP. The evaluation was performed on the 600 instances of the Benchmark Set using the four metrics defined in Section IV.1. The complete results are summarized in Fig. 4.
@@
Figure 5:  Comparative analysis of IPSA (blue), 3-PSA (purple), and IQOAP (red) on the 200 instances of the LLL-Challenging Set ($n=6$). The classical LLL algorithm (yellow-green) is also included for reference. The panels show (a) Success Rate (SR), (b) Approximation Ratio (AR), (c) Total Circuit Depth ($D_{\text{total}}$), and (d) Total CNOT Count ($C_{\text{total}}$). The y-axis values for (c) and (d) are in units of $10^{5}$. By the selection criteria for this instance set, the SR for the LLL algorithm is zero, but its average AR is high at 0.978.
The advantage of IPSA extends to the quality of approximate solutions. The AAR for IPSA remained close to 1, with a value of 0.9835 for $n=6$. While IQOAP also provides approximate solutions, its AAR values of 0.641, 0.396, and 0.294 were significantly lower than those of IPSA. The PSA variants yielded the lowest AAR values, mainly in the 0.3 to 0.5 range. This result suggests that while iterative algorithms like IPSA and IQOAP find better quality solutions than the non-iterative PSA, the partitioning strategy within IPSA provides a crucial further enhancement.
The resource requirements present a more varied picture. For $n=4$ and 5, the $D_{\text{total}}$ of IPSA was comparable to or lower than that of IQOAP and the higher-qubit $k$-PSA variants. For the largest dimension $n=6$, the $D_{\text{total}}$ for IPSA at $0.321\times 10^{6}$ was higher than for 3-PSA and IQOAP. IQOAP’s lower depth, however, is mainly attributable to its termination after a fixed 50 iterations, a number insufficient to reliably solve the problem at this scale. In terms of the $C_{\text{total}}$, IPSA consistently required fewer gates than IQOAP but more than the PSA variants. This increased resource consumption is a trade-off for the substantial gains in success rate and solution quality.
In summary, the results indicate that both PSA and IQOAP find the shortest vector with low probability, making them unreliable for this task. This low reliability stems from PSA’s tendency to converge on local minima and IQOAP’s propensity for ineffective iterative refinements that fail to improve solution quality. Their limited success at a modest dimension of $n=6$ suggests poor scalability for higher-dimensional problems. In contrast, IPSA achieves the correct solution with high probability. Its integration of iterative refinement and a novel partitioning strategy demonstrates a more effective path to achieving high-quality solutions for the SVP.
IV.4  Effectiveness on Instances with Close Suboptimal Solutions
@@
In summary, the effectiveness of IPSA on the LLL-Challenging Set highlights the efficacy of its 1-tailed search spaces strategy. This design feature equips the algorithm to perform well in scenarios with substantial suboptimal solution interference, a valuable capability for solving the SVP and other challenging combinatorial optimization problems.
V  Conclusion
In this work, we proposed the IPSA, a “second-generation” iterative VQA-based SVP solver that inherits the ideas of the PSA and the IQOAP framework. It is designed to overcome their fundamental drawbacks, such as ineffective iterations, oversized search spaces, and low circuit efficiency. Its core features, the 1-tailed search space and a stack-based iterative framework, are the key components for achieving this goal. Although PSA and IQOAP may not be the best approach for solving SVP with VQA today, the improved and refined IPSA presented in this work stands as a competitive alternative by mitigating the known drawbacks systematically, offering its own unique merits. Furthermore, IPSA can be employed as a subroutine within algorithms like block Korkin-Zolotarev or be applied to solving problems such as learning with errors, providing an efficient and reliable core component for these complex tasks.
Acknowledgements.
This work is supported by National Natural Science Foundation of China (Grant Nos. 62372048, 62272056, 62371069), and National Key Laboratory of Security Communication Foundation (2025, 6142103042503).

### v2

@@ C.2 Results on the Benchmark Set
We now compare the effectiveness of IPSA against the various $k$-PSA and IQOAP. The complete results are summarized in Fig. 6.
Figure 6:  Comparative analysis of IPSA (blue) and several existing algorithms, including the $k$-PSA variants (3-PSA in purple, 4-PSA in orange, and 5-PSA in green) and IQOAP (red). The comparison is across dimensions $n\in\{4,5,6\}$ on the Benchmark Set. The panels show (a) Success Rate (SR), (b) Approximation Ratio (AR), (c) Total Circuit Depth ($D_{\text{total}}$), and (d) Total CZ Count ($C_{\text{total}}$). The y-axis values for (c) and (d) are in units of $10^{5}$. In panels (a) and (b), the brown bars labeled “Initial” represent the metrics of the initial basis. For the data points, the error bars indicate the interquartile range (25th to 75th percentile), and the circular markers denote the median values.
@@
We note that the SR for IQOAP in these noiseless simulations is lower than that observed in the real hardware experiments presented in the main text. This discrepancy arises from the increased difficulty of the instances used in this appendix. As detailed in Appendix C.1, these simulations employed unimodular matrices with larger entries to create more complex lattice geometries, whereas the hardware experiments in the main text used instances with a restricted coefficient range. Even in this noiseless environment, the performance of IQOAP and $k$-PSA degrades significantly on these more challenging instances, whereas IPSA maintains a high SR. This outcome suggests that IQOAP’s fixed iteration limit is insufficient for these more complex instances. Although increasing the number of iterations might marginally improve the success rate, this reliance on a fixed iteration count appears to be a structural limitation that hinders scalability, as evidenced by the rapid performance decline at $n=5$ and $n=6$.
The resource requirements present a more varied picture. For $n=4$ and 5, the $D_{\text{total}}$ of IPSA was comparable to or lower than that of IQOAP and the higher-qubit $k$-PSA variants. For the largest dimension $n=6$, the $D_{\text{total}}$ for IPSA at $0.321\times 10^{6}$ was higher than for 3-PSA and IQOAP. IQOAP’s lower depth, however, is mainly attributable to its termination after a fixed 50 iterations, a number insufficient to reliably solve the problem at a little larger scale. In terms of the $C_{\text{total}}$, IPSA consistently required fewer CZ gates than IQOAP but more than the PSA variants. This increased resource consumption is a trade-off for the substantial gains in success rate and solution quality.
Crucially, it is important to note that the $D_{\text{total}}$ of IQOAP is comparable to that of IPSA in these numerical simulations because the results reflect only the logical circuit depth. This metric excludes the heavy “transpile tax” incurred by limited connectivity on real quantum computers, which explains the depth parity observed in this ideal simulation environment.
The results indicate that both PSA and IQOAP find the shortest vector with low probability, making them unreliable for this task. This low reliability stems from PSA’s tendency to converge on local minima and IQOAP’s propensity for ineffective iterative refinements that fail to improve solution quality. Their limited success at a modest dimension of $n=6$ suggests poor scalability for higher-dimensional problems. In contrast, IPSA achieves the correct solution with high probability. Its integration of iterative refinement and a novel partitioning strategy demonstrates a more effective path to achieving high-quality solutions for the SVP.
C.3 Results on the LLL-Challenging Set
@@
Figure 7:  Comparative analysis of IPSA (blue), 3-PSA (purple), and IQOAP (red) on the 200 instances of the LLL-Challenging Set ($n=6$). The classical LLL algorithm (yellow-green) is also included for reference. The panels show (a) Success Rate (SR), (b) Approximation Ratio (AR), (c) Total Circuit Depth ($D_{\text{total}}$), and (d) Total CZ Count ($C_{\text{total}}$). The y-axis values for (c) and (d) are in units of $10^{5}$. By the selection criteria for this instance set, the SR for the LLL algorithm is zero, but its average AR is high at 0.978.
@@
The effectiveness of IPSA on the LLL-Challenging Set highlights the efficacy of its 1-tailed search spaces strategy. This design feature equips the algorithm to perform well in scenarios with substantial suboptimal solution interference, a valuable capability for solving the SVP and other challenging combinatorial optimization problems.
C.4 Summary of Numerical Findings
The extensive numerical simulations reinforce and extend the findings from the hardware experiments. They confirm that IPSA’s advantages in success rate and solution quality are not limited to small-scale instances but scale to harder, higher-dimensional problems effectively. The comparison with multiple PSA variants demonstrates that IPSA’s iterative, stack-managed architecture provides a fundamental algorithmic advantage over static approaches, regardless of their specific qubit resource allocation.